OpenStudy (anonymous):

Solve for x: 5 2 2 ------ - ---- = ---- x^2 - 1 x x + 1 The answers the give me are A. x= 1/4 and x = 2 B. x= -1/4 and x = - 2 C. x= 1/4 and x= -2 D. x= -1/4 and x = 2 ?

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

How do you figure that out? If you don't mind explaining? @prathameshtambe007

OpenStudy (anonymous):

yup I'll explain. Just its slightly lengthy. have you done any work? got any answer or reached upto any stage and stuck? if its so just tell me where and if not I'll write all the steps.

OpenStudy (anonymous):

No, i wasn't' sure how to go about this? I read the lesson,but i didnt exactly understand it.

OpenStudy (anonymous):

5/(x^2 - 1) - 2/x = 2/(x+1) 5/(x^2-1) = 2/x + 2/ (x+1) 5/ (x+1)*(x-1) = (2*(x+1) + 2*x )/ x* (x+1) 5 / (x-1) = (2x + 2x + 2) / x 5/ (x-1) = 4x+2 / x 5x = (x-1) * (4x+2) 5x = 4x^2 + 2x - 4x - 2 4x^2 -7x -2 = 0 4x^2 - 8x +x -2 = 0 4x * (x-2) + (x-2) = 0 (x-2) * (4x+1) = 0 x-2 = 0 or 4x+1=0 x=2 or x= -1/4

OpenStudy (anonymous):

@brazukinhaa hope you got the answer

OpenStudy (anonymous):

you want to make the denominator equal in order to be able to add or subtract them

OpenStudy (anonymous):

Okaay, i had to write it down but i get itt. Thanks so much(: Both of you!

OpenStudy (anonymous):

most welcome buddy let me know anytime if you stuck at any step.

OpenStudy (anonymous):

Thank you, i appreciate itt(: !!